# Inferring Belief States in Partially-Observable Human-Robot Teams

### v1

@@ Abstract
We investigate the real-time estimation of human situation awareness using observations from a robot teammate with limited visibility. In human factors and human-autonomy teaming, it is recognized that individuals navigate their environments using an internal mental simulation, or mental model. The mental model informs cognitive processes including situation awareness, contextual reasoning, and task planning. In teaming domains, the mental model includes a team model of each teammate’s beliefs and capabilities, enabling fluent teamwork without the need for explicit communication. However, little work has applied team models to human-robot teaming. We compare the performance of two current methods at estimating user situation awareness over varying visibility conditions. Our results indicate that the methods are largely resilient to low-visibility conditions in our domain, however opportunities exist to improve their overall performance.
@@ I Introduction
The human factors community recognizes that when individuals conduct taskwork, they construct an internal simulation termed a mental model. The mental model broadly informs the individual’s abstract understanding of the environment, such as contextual information, the status of the task at hand, contingencies, and actionable plans [3, 17].
@@
Our results indicate that both models were successful at inferring user situation awareness. Moreover, both models maintained their performance under varying visibility conditions. The LLM performed on par with the hand-crafted logical predicates model, indicating that the LLM can be leveraged to create rules for logical predicates models, or encapsulate the process entirely.
All data and source code for this paper is available at https://jackkolb.com/tmm-hri.
@@ II Related Works
The HRI community has primarily utilized mental models with task-dependent factors aligned to the three levels of situation awareness. For example, in a process-driven task (e.g., manufacturing, housework, spaceflight), the mental model may contain a representation of the overall procedure (Level 1), where in the procedure the environment is currently (Level 2), and what that implies for the immediate future (Level 3). Contextual information is derived from the world state, which supports higher-level reasoning about the future of the environment and task. In addition, mental models can draw upon prior experiences to form a stronger model of the risks, failure modes, and practical effects of different actions and task strategies [22].
@@ II-A Logical Predicates Representation
The most prevalent representation of mental models in the HRI literature uses logical predicates, or sets of known logical truths [7, 18, 10, 19]. Each predicate is an observed feature of the environment state taken as an absolute truth, e.g., the location of an object and its properties. Complex functions are hand-crafted to derive contextual information or model how the environment will change over time. When predicting the mental model of a teammate, a robot agent can represent the teammate’s mental model using the same logical predicates framework, and filter the observational inputs by the workspace region visible to the teammate.
The logical predicates model is attractive for environments with low observational uncertainty. When there is little room for the human teammate to misunderstand their observations, and the robot has full observability, the teammate’s mental model is a subset of the robot’s mental model. With these assumptions, a rigorously-designed logical predicates model can capture many if not all the important aspects of the environment and answer a wide range of inference questions.
However, the robot’s mental model can be misled when it does not contain all information available to the teammate’s mental model, for example, in partially observable conditions. As the logical predicates model is fact-based and cannot model uncertainty, hard-coding reasoning rules to resolve conflicting information can negatively impact downstream planning and reasoning.
@@
In this work we assess whether an LLM can match a logical predicates model at inference over a scene graph, challenging the need for hand-crafted inference rules.
@@ II-B Markov Decision Process Representation
II-C Fuzzy Mental Model Finite State Machine (FMMFSM)
Bolton et al. recently proposed FMMFSM, a model inspired by fuzzy logic concepts that uses membership functions to estimate a teammate’s belief state [2]. FMMFSM addresses the uncertainty problem of the logical predicates model, and uses an object-based scene graph. While the logical predicates model translates directly from observations to an updated mental model, FMMFSM introduces three intermediate functions:
  • $\alpha$
    maps possible environment events to the probability that each event is perceived by the teammate (addressing the observability limitation of logical predicates).
  • $\phi$
    maps perceived environment events to the probability that the event was correctly perceived.
  • $\delta$
    maps perceived environment events to the probability that the event changes each aspect of the system state.
The function implementations are arbitrary and not limited to traditional hand-crafted fuzzy membership functions.
While FMMFSM addresses key shortcomings of the logical predicates and POMDP models, it solely targets Level 1 situation awareness, or the construction of the scene graph. Predictions of contextual or future information would need to be derived from its output. As this work focuses on inference over a scene graph, instead of the scene graph itself, we chose not to utilize FMMFSM in this work. However, FMMFSM presents a compelling model that warrants evaluation in real-world human-robot teaming domains.
@@ III Problem Formulation
Our objective is to map a set of observable environment states $O$ to a set of belief states $\{\beta^{true},\beta^{robot},\beta^{pred}\}$ over a time series. The belief states represent task-relevant world state and contextual information about the environment for an oracle ($\beta^{true}$), a robot ($\beta^{robot}$), and a human user ($\beta^{pred}$). $\beta^{true}$ is the ground truth representation of the environment, and is updated from the full environment state $O_{t}$ at each timestep $t$. $\beta^{robot}$ uses the subset of $O_{t}$ that is observable by the robot within an observable region, $R^{robot}$. Lastly, $\beta^{pred}$ is a predicted belief state of the human user from the robot’s perspective, updated using $\beta^{robot}$ and the user’s observable region $R^{user}$.
@@
    A fully observable logical predicates baseline ($\mathcal{B}^{LP}$) will result in responses to situation awareness queries that moderately agree with user responses.
@@
    The performance of the logical predicates baseline ($\mathcal{B}^{LP}$) at inferring user situation awareness will sharply degrade in partially observable conditions.
@@
    A large language model ($\mathcal{B}^{LP+LLM}$) will match the logical predicates baseline ($\mathcal{B}^{LP}$) across fully observable and partially observable environments.
@@ IV Methods
In Overcooked-AI, agents in a human-robot team perform a sequence of subtasks to cook soups. Agents first move three ingredients (onions or tomatoes) from countertops to a pot. The pot then cooks for $10$ seconds, and once complete, an agent ‘plates’ the soup using a dish. Lastly, the plated soup is brought to a designated serving station, where it leaves the gameboard. The process repeats until all ingredients have been used or $90$ seconds have elapsed. Users are told to focus on completing all possible soups, but are not evaluated on their performance.
@@ IV-B Object Permanence in the Scene Graph
The final pass addresses transformed objects. These are observed objects that have no fit to a known object (i.e., a newly-made soup), or known objects that have no observed match (i.e., a plated soup that was delivered). Due to the partially observable environment, it can be challenging to determine which objects transformed into other objects, however many cases can be inferred by tracking a soup’s ingredients. For example, a new “tomato, onion, and tomato soup” object implies that two known tomatoes and one known onion no longer exist on the gameboard. As each agent’s belief state is initialized with the true gameboard state, and we can assume no ingredients are added during gameplay, this pass can reasonably identify which ingredients were used to create a soup and when a soup has likely been served.
Due to the freedom users have to select ingredients and complete the task, no object permanence system can be perfect in all partially observable environments.
@@ IV-C Inferring Context and Situation Awareness
The value of modeling a teammate’s belief state lies in applying the belief state to correctly infer additional contextual information and answer questions. We use the SAGAT method [8] to obtain the user’s real belief state, $\beta^{human}$, by periodically pausing the game and asking the user questions about the environment. Up to two questions are asked every $30$ seconds. We include questions about the first two levels of situation awareness – world state understanding and contextual information. All situation awareness questions were multiple choice and randomly chosen from a set question bank.
@@
For the logical predicates model, $\mathcal{B}^{LP}$, hand-crafted rules use the current scene graph to derive responses to each situation awareness question. For example, to answer the contextual information question “How many more soups can be made/delivered?”, $\mathcal{B}^{LP}$ calculates the maximum number of remaining soups using the ingredients and plates known to $\beta^{pred}$. Consequently, the logical predicates model is unable to answer questions that are not hand-crafted beforehand.
$\mathcal{B}^{LP+LLM}$ uses $\mathcal{B}^{LP}$ to construct the predicted scene graph component of $\beta^{pred}$, and an LLM (GPT4 [1]) to reason upon $\beta^{pred}$ to answer situation awareness questions. The model is prompted with a description of the game and the current scene graph representation, as well as the query “You are A1 […] Please answer the question using only one of responses below […] What is your answer?” To fairly compare $\mathcal{B}^{LP}$ and $\mathcal{B}^{LP+LLM}$, the models do not consider prior world states nor data from other participants, although exploring alternative methods of constructing the inferred scene graph is an opportunity for future work. The primary benefit of using an LLM for this task is the LLM’s ability to answer open vocabulary questions about the scene, which could supercede a logical predicates model. Additionally, an LLM can distill its reasoning into a logical format compatible with a $\mathcal{B}^{LP}$ model for efficient repeat queries.
@@ IV-D Scoring and Evaluation
Belief states are scored by comparing the responses of the belief states to situation awareness questions. Correct responses are awarded one point, and incorrect responses zero. For questions based on spatial reasoning, we awarded partial credit for near-correct answers. For example, for a question asking “Where is the closest tomato?”, responding “Center” when the correct answer was one tile off-center scored $0.5$. The scoring function is defined as:
$Score_{u,l}=\frac{\sum\limits_{q,t\in N}(score_{q}(\beta^{A}_{t},\beta^{B}_{t})}{|N|}$ (3)
@@ V User Study Design
Participants completed two practice runs and then four trials with our experiment domain. Each trial used a different kitchen layout. The game’s state was recorded at each frame ($10$hz), as well as responses to situation awareness questions.
@@
Figure 4: Violin plot of the score distributions of user responses to situation awareness questions ($\beta^{human}$) with respect to the ground truth oracle ($\beta^{true}$). The broad variances show the task is appropriately difficult.
@@ VI Results & Discussion
A total of $42$ participants completed the study ($50\%$ female, ages $22-74$). We excluded $11$ participants who either failed to complete the study or failed to outperform random selection of the situation awareness responses, resulting in $31$ participants included in our evaluation.
@@ VI-A User Performance
Notably, few users scored perfectly (scores of $100\%$), and no users scored consistently well across all layouts. This demonstrates that users were generally unable to fully track the environment, motivating the potential benefits of predicting user belief states in this partially observable domain.
VI-B Fully Observable Robot Belief State
Hypothesis H1 evaluates the logical predicates baseline in its ideal conditions. We are interested in how accurately a robot with full observability can predict the situation awareness responses of a human user with partial observability. Unlike our evaluation of the User Performance, wherein the human user was scored against the ground truth oracle, in this case the robot does not assume that the user has any knowledge of the environment outside of what the robot knows the user has explicitly seen.
We find that the scores from the predicted user belief state ($\beta^{pred}$) with respect to the actual user responses ($\beta^{human})$ are almost identical to Fig. 4. The finding indicates that there were virtually no instances in our study where a model assuming perfect user perception could have had a false belief, i.e., environment changes that would have impacted the situation awareness responses were quickly observed by users, although seemingly not always recalled.
Additionally, Fig. 2 shows that the full observability model’s performance averaged between $55\%$ and $75\%$. This suggests that users had nuanced perception, comprehension, or memory that was not captured by $\mathcal{B}^{LP}$. Therefore, opportunities exist for alternative models to utilize human factors that can contribute to a user’s belief state.
@@ VI-B Fully Observable Robot Belief State
Figure 5: Line plot of the mean performance of $\beta^{pred}$ using $\mathcal{B}^{LP}$ with respect to the user responses, across various visibility parameters $R^{robot}$. While performance declines between the three visibility types, there is not a meaningful decline between visibility conditions.
@@ VI-C Partial Robot Observability
Fig. 5 shows the performance of $\mathcal{B}^{LP}$ across $14$ visibility conditions. Counter to our hypothesis H2, the performance of $\mathcal{B}^{LP}$ only slightly decreased between high visibility and low visibility conditions. We believe this due to our logical predicates implementation accurately resolving conflicting beliefs during the task. The findings support using $\mathcal{B}^{LP}$ as a baseline for more complex partially observable domains.
@@
Figure 6: Chart of the mean performances of $\beta^{pred}$ constructed using the $\mathcal{B}^{LP}$ and $\mathcal{B}^{LP+LLM}$ models over two observability conditions. Error bars show the variance. $\mathcal{B}^{LP+LLM}$ has comparable means and variances to $\mathcal{B}^{LP}$ when scored against user responses, indicating that $\mathcal{B}^{LP+LLM}$ is able to infer an inference ruleset similar to the hand-crafted developed for $\mathcal{B}^{LP}$. Higher scores indicate superior performance.
@@ VI-D Large Language Model
Hypothesis H3 explores whether an LLM can predict and reason about the user’s belief state as a zero-shot system. The open vocabulary reasoning capability of LLMs has no equivalent for logical predicates, and would be valuable for broadly applying predicted belief states. As a first step, we are interested in whether an LLM can match the performance of the logical predicates model ($\mathcal{B}^{LP}$).
Fig. 6 shows the performance of $\mathcal{B}^{LP}$ and $\mathcal{B}^{LP+LLM}$ at predicting user responses to situation awareness questions. The two models were presented with the same scene graph and prompted with which answer the user would choose, and we found no meaningful difference in mean nor variance between the two models. The findings support the use of LLMs for predicting user belief states, and as a baseline upon which to compare future models.

### v2

@@ Abstract
We investigate the real-time estimation of human situation awareness using observations from a robot teammate with limited visibility. In human factors and human-autonomy teaming, it is recognized that individuals navigate their environments using an internal mental simulation, or mental model. The mental model informs cognitive processes including situation awareness, contextual reasoning, and task planning. In teaming domains, the mental model includes a team model of each teammate’s beliefs and capabilities, enabling fluent teamwork without the need for explicit communication. However, little work has applied team models to human-robot teaming. In this work we compare the performance of two models, logical predicates and large language models, at estimating user situation awareness over varying visibility conditions. Our results indicate that the methods are largely resilient to low-visibility conditions in our domain, however opportunities exist to improve their overall performance.
@@ I Introduction
The human factors community recognizes that when individuals conduct taskwork, they construct an internal simulation termed a mental model. The mental model broadly informs the individual’s abstract understanding of the environment, such as contextual information, the status of the task at hand, contingencies, and actionable plans [17].
@@
Our results indicate that both models were successful at inferring user situation awareness. Moreover, both models maintained their performance under varying visibility conditions. The LLM performed on par with the handcrafted logical predicates model, indicating that the LLM can be leveraged to create rules for logical predicates models, or encapsulate the process entirely.
A project website with a demo, dataset, and source code is available at https://jackkolb.com/tmm-hri.
@@ II Related Works
The HRI community has primarily designed mental models with task-dependent factors aligned to Endsley’s three levels of situation awareness. For example, in a process-driven task (e.g., manufacturing, housework, spaceflight), the mental model would contain the task procedure (Level 1), the context of the environment state with respect to the task procedure (Level 2), and what that implies for the immediate future (Level 3). Contextual information is derived from the world state, which supports higher-level reasoning about the future of the environment and task. In addition, mental models can draw upon prior experiences to form a stronger model of the risks, failure modes, and practical effects of different actions and task strategies [22].
@@ II-A Logical Predicates Representation
The most prevalent representation of mental models in the HRI literature uses logical predicates, or sets of known logical truths [6, 18, 9, 19]. Each predicate is an observed feature of the environment state taken as an absolute truth, e.g., the location of an object and its properties. Complex functions are handcrafted to derive contextual information or model how the environment will change over time. When predicting the mental model of a teammate, a robot agent can represent the teammate’s mental model using the same logical predicates framework, and filter the observational inputs by the workspace region visible to the teammate.
The logical predicates model is attractive for environments with low observational uncertainty. When the human teammate is unlikely to misunderstand their observations, and the robot has full observability, the teammate’s mental model is a subset of the robot’s mental model. With these assumptions a logical predicates model can capture all relevant environment features and answer a wide range of inference questions.
However, the robot’s mental model can be misled when it does not contain all information available to the teammate’s mental model, for example, in partially-observable conditions. As the logical predicates model is fact-based and cannot model uncertainty, hard-coding reasoning rules to resolve conflicting information can negatively impact downstream planning and reasoning.
@@
In this work we assess whether an LLM can match a logical predicates model at inference over a scene graph, challenging the need for handcrafted inference rules.
@@ II-B Markov Decision Process Representation
II-C Scene Graphs
In practice, a belief state is built upon a representation of the world state. Prior work typically uses a semantic scene graph of key environment objects and properties, reducing the complexity of downstream reasoning and inference [10].
A scene graph is defined as a graph $G=\{V,E\}$. Graph nodes $V$ represent environment objects, with properties such as the object’s location, class, and instance-level attributes. Graph edges $E$ represent relationships between nodes, with properties such as the semantic relationship between objects (e.g., onTopOf). Nodes are not limited to environment objects and can include higher-level concepts such as rooms and structures [16]. Similarly, edges can include information about what tasks an object is useful for and their decomposition. As current representations of mental models are task-dependent, the scene graph offers a flexible structure for representing physical and abstract concepts in the scene.
In a dynamic scene graph (DSG), the scene changes with time and $G$ is updated as the agent receives new information from cameras or other sensors. In partially-observable domains – i.e., a mobile robot with a camera – maintaining a DSG is highly challenging because of object impermanence. A DSG must not only track where objects are, but resolve new, moved, and removed object instances. Consequently, no libraries exist for constructing DSGs in partially-observable environments. Instead, related works often extract the scene graph from the simulator itself or simplify the domain to avoid instance-level reasoning (e.g., no two objects of the same class and color). In this work we introduce a basic – but functional – model for maintaining a DSG.
@@ III Problem Formulation
Our objective is to map a set of observable environment states $O$ to a set of belief states $\{\beta^{true},\beta^{robot},\beta^{pred}\}$ over a time series. The belief states represent task-relevant world state and contextual information about the environment for an oracle ($\beta^{true}$), a robot ($\beta^{robot}$), and a human user ($\beta^{pred}$). $\beta^{true}$ is the ground truth of the environment, and is updated from the full environment state $O_{t}$ at each timestep $t$. $\beta^{robot}$ uses the subset of $O_{t}$ that is observable by the robot within an observable region, $R^{robot}$. Lastly, $\beta^{pred}$ is a predicted belief state of the human user from the robot’s perspective, updated using $\beta^{robot}$ and the user’s observable region $R^{user}$.
@@
    A fully-observable logical predicates baseline ($\mathcal{B}^{LP}$) will result in responses to situation awareness queries that agree with user responses.
@@
    The performance of the logical predicates baseline ($\mathcal{B}^{LP}$) at inferring user situation awareness will sharply degrade in partially-observable conditions.
@@
    A large language model ($\mathcal{B}^{LP+LLM}$) will match the logical predicates baseline ($\mathcal{B}^{LP}$) across fully-observable and partially-observable environments.
@@ IV Methods
In Overcooked-AI, agents in a human-robot team perform a sequence of subtasks to cook soups. Agents first move three ingredients (onions or tomatoes) from countertops to a pot. The pot then cooks for $10$ seconds, and once complete, an agent ‘plates’ the soup using a dish. Lastly, the plated soup is brought to a designated serving station, where it leaves the gameboard. The process repeats until all ingredients have been used or $90$ seconds have elapsed. Users are told to focus on completing all possible soups. While each user’s time to complete and quantity of completed soups are available in the dataset, we only used user performance to screen out participants who ignored the task. Participants were filtered out if they answered no situation awareness questions correctly, or failed to cook at least one soup in any of the four kitchen layouts.
@@ IV-B Object Permanence in the Scene Graph
The final pass addresses transformed objects. These are observed objects that have no fit to a known object (i.e., a newly-made soup), or known objects that have no observed match (i.e., a plated soup that was delivered). Due to the partially-observable environment, it can be challenging to determine which objects transformed into other objects, however many cases can be inferred by tracking a soup’s ingredients. For example, a new “tomato, onion, and tomato soup” object implies that two known tomatoes and one known onion no longer exist on the gameboard. As each agent’s belief state is initialized with the true gameboard state, and no ingredients are added during gameplay, this pass can reasonably identify which ingredients were used to create a soup and when a soup has likely been served.
Due to the freedom users have to select ingredients and complete the task, no object permanence system can be perfect in all partially-observable environments.
@@ IV-C Inferring Context and Situation Awareness
The value of modeling a teammate’s belief state lies in applying the belief state to correctly infer additional contextual information and answer questions. We use the SAGAT method [7] to obtain the user’s real belief state, $\beta^{human}$, by periodically pausing the game and asking the user two questions about the environment every $30$ seconds. We include questions about the first two levels of situation awareness – world state understanding and contextual information. All situation awareness questions were multiple choice and randomly chosen from a set question bank.
@@
For the logical predicates model, $\mathcal{B}^{LP}$, handcrafted rules use the current scene graph to derive responses to each situation awareness question. For example, to answer the contextual information question “How many more soups can be made/delivered?”, $\mathcal{B}^{LP}$ calculates the maximum number of remaining soups using the ingredients and plates known to $\beta^{pred}$. Consequently, the logical predicates model is unable to answer questions that are not handcrafted beforehand. The model’s specific rules are detailed on the project webpage.
$\mathcal{B}^{LP+LLM}$ uses $\mathcal{B}^{LP}$ to construct the predicted scene graph component of $\beta^{pred}$, and an LLM (GPT4 [1]) to reason upon $\beta^{pred}$ to answer situation awareness questions. The model is prompted with a description of the game and the scene graph $\beta^{pred}$, as well as the query “You are A1 […] Please answer the question using only one of responses below […] What is your answer?” To fairly compare $\mathcal{B}^{LP}$ and $\mathcal{B}^{LP+LLM}$, the models do not consider prior world states nor data from other participants, although exploring alternative methods of constructing the inferred scene graph is an opportunity for future work. The primary benefit of using an LLM for this task is the LLM’s ability to answer open vocabulary questions about the scene, which removes the human effort needed to define the rules of an LP model (albeit at a high computational cost). Additionally, an LLM can distill its reasoning into a logical format compatible with a $\mathcal{B}^{LP}$ model for efficient repeat queries. Our specific LLM prompt is on the project webpage.
@@ IV-D Scoring and Evaluation
Belief states are scored by comparing the model’s responses of the belief states to situation awareness questions. Correct responses are awarded one point, and incorrect responses zero. For questions based on spatial reasoning, near-correct answers received 0.5 points. For example, for a question asking “Where is the closest tomato?”, responding “Center” when the correct answer was one tile off-center scored $0.5$. Partial-credit was used because the situation awareness questions were multiple choice, so a distance metric was infeasible. The scoring function is defined as:
$Score_{u,l}=\frac{\sum\limits_{q,t\in N}score_{q}(\beta^{A}_{t},\beta^{B}_{t})}{|N|}$ (3)
@@ V User Study Design
Participants completed two practice runs and then four trials with our experiment domain. Each trial used a different kitchen layout shown in Fig. 7. The game’s state was recorded at each frame ($10$hz), as well as responses to situation awareness questions.
@@
Figure 4: The score distributions of user responses to situation awareness questions ($\beta^{human}$) with respect to the ground truth ($\beta^{true}$). The broad variances show the task is appropriately difficult. Fig. 7 shows the layouts.
@@ VI Results & Discussion
We continued recruiting participants until $30$ participants passed the minimum participation conditions. A total of $42$ participants were recruited ($50\%$ female, ages $22-74$), of whom $30$ are considered in the dataset and model evaluation.
@@ VI-A User Performance
Notably, very few users scored perfectly (scores of $100\%$), and no users scored consistently well across all layouts. This suggests that users were generally unable to fully track the environment, motivating the potential benefits of predicting user belief states in this partially-observable domain.
VI-B Fully-Observable Robot Belief State
Hypothesis H1 evaluates how accurately $\mathcal{B}^{LP}$ with full observability can predict the situation awareness responses of a human with partial observability. Unlike our evaluation of the User Performance, wherein the human user was scored against the ground truth oracle, in this case the robot does not assume that the user has any knowledge of the environment outside of what the robot knows the user has explicitly seen.
We find that the scores from the predicted user belief state ($\beta^{pred}$) with respect to the actual user responses ($\beta^{human})$ are almost identical to Fig. 4. This indicates that there were few instances in our study where a model assuming perfect user perception could have had a false belief, i.e., environment changes that would have impacted the situation awareness responses were quickly observed by users, although seemingly not always recalled.
Similarly, we find that the full observability model’s performance averaged between $55\%$ and $75\%$. This suggests that users had nuanced perception, comprehension, or memory that was not captured by $\mathcal{B}^{LP}$. Opportunities exist for alternative models to utilize human factors that can contribute to a user’s belief state.
@@ VI-B Fully-Observable Robot Belief State
Figure 5: Line plot of the performance of $\beta^{pred}$ using $\mathcal{B}^{LP}$ with respect to the user responses, across visibility parameters $R^{robot}$. Error bars indicate variance. While performance declines between the three visibility types, there is not a significant decline between visibility conditions.
@@ VI-C Partial Robot Observability
Fig. 5 shows the performance of $\mathcal{B}^{LP}$ across $14$ visibility conditions. Counter to our hypothesis H2, the performance of $\mathcal{B}^{LP}$ did not significantly decrease between high visibility and low visibility conditions (2-sample t-test, p$>$$0.1$ for all). We believe this due to our logical predicates implementation accurately resolving conflicting beliefs during the task. The findings support using $\mathcal{B}^{LP}$ as a baseline for more complex partially-observable domains.
@@
Figure 6: Mean performances of $\beta^{pred}$ constructed using $\mathcal{B}^{LP}$ and $\mathcal{B}^{LP+LLM}$ over two observability conditions. Higher scores indicate superior performance, error bars show variance. $\mathcal{B}^{LP+LLM}$ matched or exceeded $\mathcal{B}^{LP}$ when scored against user responses, indicating that $\mathcal{B}^{LP+LLM}$ is able to infer an ruleset similar to $\mathcal{B}^{LP}$’s handcrafted rules.
@@ VI-D Large Language Model
Hypothesis H3 explores whether an LLM can predict and reason about the user’s belief state as a zero-shot system. The open vocabulary reasoning of LLMs departs from the handcrafted rules of logical predicates models, adding value for flexible real-world domains. As a first step, we are interested in whether an LLM can match the performance of $\mathcal{B}^{LP}$.
Fig. 6 shows the performance of $\mathcal{B}^{LP}$ and $\mathcal{B}^{LP+LLM}$ at predicting user responses to situation awareness questions. The models were presented with the same scene graph and prompted with which answer the user would choose. Surprisingly, $\mathcal{B}^{LP+LLM}$ outperformed $\mathcal{B}^{LP}$ in full observability (2-sample t-test, p=$0.09$), while the models had near-identical performances given partial-visibility (O4). This suggests that LLMs are capable of inferring similar rules to handcrafted models in zero-shot. The findings support using LLMs to predict user belief states, and as a baseline in future work.